OpenStudy (anonymous):

For the sequence given by an (a sub n) = 4n+5, find the first 4 terms of the sequence by substituting n=1,2,3,4 into the equation.

OpenStudy (shamil98):

me to the rescue lmao. a(1) = 4 + 5 = 9 a(2) = 8 + 5 = 13 a(3) = 12 + 5 = 17 a(4) = 16 + 5 = 21

OpenStudy (anonymous):

9, 13, 17, 21?

OpenStudy (anonymous):

yay!! :)

OpenStudy (shamil98):

yes.

OpenStudy (anonymous):

now how would i identify the 1st term and the 10th term of the sequence?

OpenStudy (anonymous):

would the first term be 9 and the last term is 4(10)+5?

OpenStudy (shamil98):

yeah.

OpenStudy (anonymous):

so 9 & 45 :D

OpenStudy (shamil98):

mhm.

OpenStudy (anonymous):

how would i use the first and tenth term to find the sum of the first tenth terms?

OpenStudy (shamil98):

S(10) = 1/2(45+9)(10) S(10) = 22(10) S(10)= 220

OpenStudy (shamil98):

Understand what i did?

OpenStudy (anonymous):

well with the 45+9, what happened to the 54?

OpenStudy (shamil98):

the sum of the 10th term = 1/2(if a(1) + if a(10)) x 10 45+9 = 54 1/2(54) = 22

OpenStudy (anonymous):

ahh i see now! thank you so much :)

OpenStudy (shamil98):

no problem.